Plz give the detailed integration of the following expression to find v(especially after putting the limits): dv/6-3v = dt

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Bhavya
 
dv/ 6 -3v. = dt
ln(6-3v)/ (-3) = t + c
ln(6-3v) = -3t -3c = -3t + k
6-3v = e^ (-3t+k)
3v = 6 - e^(-3t+k)
 
Regards
Arun (askIITians forum expert)